OpenStudy (anonymous):

Translate the word phrase into a math expression. six less than eight times a number A. 6n – 8 B. 6 – 8n C. 8n – 6 D. 6 < 8n I think the answer is D?

OpenStudy (anonymous):

@silverr @dtan5457

OpenStudy (anonymous):

8n-6

OpenStudy (anonymous):

@sweetness221 Thank you! How'd you get it so fast?!

OpenStudy (anonymous):

@sweetness221 could you help me with 5 or so more?:D or 4..

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Which property justifies this statement? If x + y = z and z = x + 4, then x + y = x + 4. A. symmetric B. associative C. distributive D. transitive @sweetness221

OpenStudy (anonymous):

@chelle._cloud

OpenStudy (anonymous):

idk on this one

OpenStudy (anonymous):

Which situation could be solved using this expression? 4k + 2 A. Jason read four more than twice the number of books that Marty read. How many books did Marty read? B. Jason read half as many books as Marty read. How many books did Jason and Marty both read? C. Jason read two more than four times the number of books that Marty read. How many books did Jason read? D. Jason read four times as many books as Marty. Marty read two books. How many books did they read together? @sweetness221

OpenStudy (anonymous):

@RedoHawk

OpenStudy (anonymous):

@silverr

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@Godlovesme

OpenStudy (anonymous):

@silverr

OpenStudy (godlovesme):

it is adding two which means "two more than" in math case so, we can eliminate A and what the expression directly saying is that someone is reading 4 books and the other person reads 2 more

OpenStudy (anonymous):

one second

OpenStudy (godlovesme):

does that sound confusing? :P @PizzaLover123

OpenStudy (anonymous):

@Godlovesme Just a tiny bit LOL

OpenStudy (godlovesme):

lol i knew it i confuse ppls :P what i meant is let's say martin reads 4 books which will be 4b and jason reads 4b and 2 additional books so what's ur guess ?

OpenStudy (anonymous):

@Godlovesme I see I think its C!

OpenStudy (godlovesme):

awesome u r correct :) and i spelled marty wrong >.? loll

OpenStudy (anonymous):

Thanks so much could you help me with 2 or 3 more?! And Marty Lol Martin is better!

OpenStudy (godlovesme):

ikr ya i spelled it wrong on purpose ;) and ya sure open a new question though

OpenStudy (anonymous):

@Godlovesme LOL and exactly how do you do that im new to this XD

OpenStudy (godlovesme):

click the close icon thingy

OpenStudy (anonymous):

Okay LOL